OpenStudy (anonymous):

A health club currently charges its 1500 members monthly membership dues of $41. The board decides to increase the amount of dues charged. The marketing dept. shows that each increase of $4 will result in the loss of 4 clients. How much should the club charge to optimize the revenue from monthly dues?

OpenStudy (anonymous):

Oops, each increase of $1 will result in the loss of 4 members

OpenStudy (anonymous):

this is approximately the same as the other one.

OpenStudy (anonymous):

did you attempt this at all?

OpenStudy (anonymous):

I got $51.25 but it says that isn't right

OpenStudy (anonymous):

lets see. they will make \[(41+x)(1500-4x)\] yes. where x is the dollar increase

OpenStudy (anonymous):

otherwise known as \[-4 x^2+1336 x+61500\]

OpenStudy (anonymous):

a parabola facing down, max is at vertex, compute \[\frac{-b}{2a}\] to get your answer

OpenStudy (anonymous):

\[\frac{1336}{8}\] dang i get $167 maybe i made a mistake

OpenStudy (anonymous):

increase of one dollar gives decrease of 4 members yes? so if you start at 41 and increase by x dollars you get 41+x dollars per member. decrease for every dollar is 4 members. so if you have 1500 members and increase by x dollars you get 1500-4x members. this equation looks good to me

OpenStudy (anonymous):

Thank you for trying :-)

OpenStudy (anonymous):

maybe i multiplied wrong hold on.

OpenStudy (anonymous):

not it is right. maybe the answer really is $167

OpenStudy (anonymous):

can you check?

OpenStudy (anonymous):

Nope, not it.

OpenStudy (anonymous):

satellite's answer looks right to me.

OpenStudy (anonymous):

Sorry, I tried it and it wasn't correct.

OpenStudy (anonymous):

i keep trying and it keeps looking right. are you sure you have the problem written correctly?

OpenStudy (anonymous):

A health club currently charges its 1500 members monthly membership dues of $41. The board decides to increase the amount of dues charged. The marketing dept. shows that each increase of $1 will result in the loss of 4 clients. How much should the club charge to optimize the revenue from monthly dues?

OpenStudy (anonymous):

hold on! satellite, did you get x = 167?

OpenStudy (anonymous):

after you differentiated?

OpenStudy (anonymous):

k i am holding

OpenStudy (anonymous):

this is not a calc problem. vertex is \[-\frac{b}{2a}\]

OpenStudy (anonymous):

Yep, double checked the problem and you wrote it correctly satellite73.

OpenStudy (anonymous):

\[\frac{1336}{8}=167\]

OpenStudy (anonymous):

it is the same.

OpenStudy (anonymous):

then i say $167 and i am sticking with that

OpenStudy (anonymous):

Ugh.

OpenStudy (anonymous):

no, 167 is the amount they charge on top of 41 to get maximum profit. so they charge 167+41

OpenStudy (anonymous):

oh for the love of peter. ho ho ho

OpenStudy (anonymous):

all together now! facepalm!

OpenStudy (anonymous):

sorry.

OpenStudy (anonymous):

Hooray! That's it!

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

Thank you both!!

OpenStudy (anonymous):

you are welcome :)

OpenStudy (anonymous):

also, satellite, if you differentiate a down facing parabola and equate it to 0, you will get the maxima which happens to be the vertex of the parabola. so yes, this is indeed a calculus problem. in fact, that is why the vertex of the parabola is -b/2a try it out with this problem and you will see what i'm saying.

OpenStudy (anonymous):

y=ax^2+bx+c dy/dx=2ax+b 0=2ax+b -b=2ax x=-b/2a solution to what the person with the hard name issaying

OpenStudy (anonymous):

dhatraditya i mean

OpenStudy (anonymous):

lol, yes